OpenStudy (anonymous):

Write the slope-intercept form of the line which passes through (4, 2) with a slope of 1/2. how do i solve this?

OpenStudy (lstrasmann):

use point slope form.. \[y - y_{1} = m(x - x_{1})\]

OpenStudy (lstrasmann):

then solve for y

OpenStudy (anonymous):

use slope point slope form y-y1=m(x-x1) m is slope

OpenStudy (anonymous):

wait its noym= y1-y/x-x1

OpenStudy (kymber):

Slope-intercept form is y = mx + b All you do is plug in the slope and the point given in the problem and then figure out the y-intercept. (b) m is the slope and x and y is the point given. (4, 2) (x, y) So plug these into the formula and solve for b. Once you find b, then you plug it back into the formula along with the slope and you're done.

OpenStudy (anonymous):

2=m4+b?

OpenStudy (kymber):

Yes, but put in the slope as m. Then you can solve for b.

OpenStudy (anonymous):

where does trthe1/2 go?

OpenStudy (anonymous):

Slope Intercept form: y = mx + b y axis = 2 so *2 = mx + b* x axis is 4 so * 2 = m(4) + b m = slope sp *2 = 1/2(4) + b now multiply 1/2 * 4 to get 2 now we have 2 = 2 + b subtract 2 from each side b = 0

OpenStudy (kymber):

1/2 is the slope. Plug it in for m.

OpenStudy (anonymous):

ok 2=1/2(4)+b anf i try to find b right?

OpenStudy (anonymous):

1. 2 = 1/2(4) + b 2. 2 = 2 + b 3. subtract 2 from each side to get b = 0

OpenStudy (kymber):

Thanks for just giving her the answer, Storm. She was about to do it on her own.

OpenStudy (anonymous):

the thing isthats nit one of the answers y = 1/2 x + 2 y = 1/2 x y = 1/2 x + 4

OpenStudy (anonymous):

your welcome

OpenStudy (kymber):

Since b is 0, you just don't need to have it there.

OpenStudy (kymber):

So you plug the slope back into the original equation once your have b. But in this case, you just plug the slope into the equation and ignore b.

OpenStudy (anonymous):

so its the second one?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok thanks peoples!

OpenStudy (kymber):

Yes. \(\Large\ddot\smile\)

OpenStudy (anonymous):

y-2=1/2(x-4) y-2=1/2x-2 add 2 on both sides y=1/2x here y intercept is 0